# Xiaoyue Wang  Jianyou Wang11footnotemark: 1  Weili Cao11footnotemark: 1  Kaicheng Wang11footnotemark: 1 Ramamohan Paturi  Leon Bergen22footnotemark: 2

### v1

@@ 3 A Benchmark for LLM-based IR Systems
Num. of Triplets
Num. of Queries
Num. of Docs
Relevant Doc. per Query
Query Length
Doc Length
Lexical Overlap
@@
DORIS-MAE
AI/ML
Retrieve relevant abstracts for the research question.
Multi-level
3k
60
100
18.2
166.7
199.5
0.147
ArguAna
Debate
Retrieve the most effective argument that counters the query.
Binary
19k
100
100
1
213.0
212.3
0.139
WhatsThatBook
Literature
Find the book from user’s ambiguous descriptions.
Binary
3k
100
100
1
173.6
219.5
0.049
Clinical-Trial
BioMed
Match patient records with suitable clinical trials.
3-level
1k
50
100
20.9
89.7
197.9
0.048
RELIC
Literature
Recover the masked quotation in a literary analysis.
Binary
68k
100
100
1
198.7
96.3
0.031
BIRCO Avg.
20k
51.3
100
8.4
168.3
185.1
0.083
BEIR Avg.
3710.9
2417k
43.5
9.8
124.9
0.384
@@
Max possible 100 100 Max possible 100 94.3 Max possible 100 100
ANCEFirstP 33.0 20.0 ANCEFirstP 81.9 49.6 E5-L-v2 50.0 38.52
SimLM 41.1 18.1 SimLM 85.2 44.6 RankLLaMA 56.6 47.42
SPLADE-v2 36.8 17.7 BM25 59.1 33.5 TART 44.8 39.48
@@ 3.3.2 Reducing contamination
MS MARCO 76 18 6
TREC-COVID 47 2 1
NQ 89 7 4
HotpotQA 60 20 20
FiQA 73 19 8
@@
DBPedia 78 12 10
FEVER 99 0 1
Climate-FEVER 81 18 1
Touché-2020 47 1 0
SciFact 71 25 4
Table 3: Data contamination in BEIR. GPT4 answers the query without access to any candidate passages. C0,1,2 corresponds to Category 0,1,2. C0 means contamination. Numbers are %.
@@ 3.7 Hard Negatives in BIRCO
Figure 3: Jaccard similarity between pairs of datasets from BIRCO
@@ 3.8 Candidate Pool Construction
We construct candidate pools for each query using the lexicon-matching algorithm BM25 Trotman et al. (2014) and state-of-the-art text embedding model ada-002. Each query has a candidate pool of 50 passages. (The ground-truth passage is inserted when necessary.) As shown in Appendix C and Table 9, the difficulty is comparable to the original datasets for four of five tasks, and remains challenging for the fifth.
@@ 5 LLM Benchmark Results
nDCG@10 Recall@5 nDCG@10 Recall@5 nDCG@10 Recall@5 nDCG@10 Recall@5 nDCG@10 Recall@5
Avg. nDCG@10
Maximum Possible 100 45.6 100 100 100 100 100 37.91 100 100
100
Random 49.9 4.5 9.1 10.0 9.1 10.0 22.3 7.6 9.0 9.9
19.9
@@
TART 64.9 10.2 47.6 63.7 42.7 46.1 32.8 10.8 9.4 7.9
39.5
TART+O 49.6 4.9 23.1 28.0 17.9 20.7 23.6 7.7 14.3 18.3
25.7
MonoT5 66.9 12.4 45.9 52.0 50.8 59.0 33.2 14.2 13.9 11.9
42.1
@@
RankStripedHyena 71.7 14.6 40.7 59.8 30.4 37.9 30.5 10.6 10.6 14.9
36.8
Rank+OStripedHyena 70.6 15.5 41.4 48.9 34.2 40.9 28.7 10.1 11.3 14.9
37.2
RankLlama2-7b 71.0 13.2 41.4 54.8 29.5 37.8 29.3 37.8 10.5 11.8
36.3
Rank+OLlama2-7b 70.8 14.1 41.1 54.6 26.7 34.8 28.8 9.7 10.6 10.8
35.9
RankLlama2-13b 71.2 13.0 40.1 53.9 26.5 36.8 29.6 10.6 10.7 14.9
35.6
Rank+OLlama2-13b 70.5 14.2 39.1 51.7 28.5 35.8 28.4 10.0 8.9 12.8
35.0
RankLlama2-70b 70.7 15.1 39.0 49.8 27.6 36.8 30.2 10.3 10.5 13.8
35.6
Rank+OLlama2-70b 70.4 14.2 41.6 51.7 27.6 36.8 28.8 9.7 10.5 13.9
35.8
RankGPT4 76.2 17.5 26.7 17.0 79.5 85.9 39.9 14.8 40.1 48.0
52.5
Rank+OGPT4 77.4 17.6 54.4 70.9 82.1 88.0 38.6 14.6 62.3 70.1
63.0
@@
Score+OStripedHyena 67.5 9.5 20.6 24.9 43.0 50.8 32.5 10.7 20.7 25.0
36.9
Reason+OStripedHyena 65.9 10.9 26.3 32.0 22.2 22.1 27.5 11.0 12.8 11.8
30.9
Subtask+OStripedHyena 74.8 15.5 14.1 20.9 44.4 49.0 32.4 11.0 24.7 29.1
38.1
Score+OLlama2-7b 71.6 14.1 27.0 40.1 30.5 34.7 29.7 10.1 10.9 13.9
33.9
Reason+OLlama2-7b 72.1 13.8 21.3 27.8 21.4 28.1 31.0 9.9 11.5 12.0
31.5
Subtask+OLlama2-7b 69.2 11.8 16.4 18.8 29.2 32.9 25.1 8.6 10.5 11.8
30.1
Score+OLlama2-13b 77.3 14.5 15.4 21.9 28.1 31.9 30.0 9.9 16.8 18.0
33.5
Reason+OLlama2-13b 60.1 7.5 31.4 43.1 18.7 27.0 24.2 7.7 9.0 8.9
28.7
Subtask+OLlama2-13b 71.3 14.1 12.4 14.9 48.7 60.9 29.4 10.7 17.6 21.0
35.9
Score+OLlama2-70b 75.1 15.1 40.3 49.8 60.8 70.0 34.9 12.1 12.2 14.9
44.7
Reason+OLlama2-70b 70.1 14.5 42.8 55.0 43.4 48.0 31.3 9.9 12.7 15.9
40.1
Subtask+OLlama2-70b 76.7 16.3 19.3 27.0 49.1 60.1 32.8 10.5 19.6 21.0
39.5
Score+OGPT4 79.9 19.3 51.6 70.0 83.3 90.9 43.4 17.2 54.1 70.1
62.5
Reason+OGPT 74.9 18.1 59.9 76.8 74.9 82.8 45.7 17.0 39.9 51.0
59.1
Subtask+OGPT4 78.5 18.5 69.5 85.9 79.6 85.9 43.5 16.5 53.2 62.9 64.9
@@ Appendix A Experiment Details
nDCG@10 R@5 nDCG@10 R@5 nDCG@10 R@5 nDCG@10 R@5 nDCG@10 R@5
Avg. nDCG@10
Maximum 100 45.6 100 100 100 100 100 37.91 100 100
100
Random 49.9 4.5 9.1 10.0 9.1 10.0 22.3 7.6 9.0 9.9
19.9
@@
RoBERTa-L 66.8 $\pm$ 1.3 12.0 $\pm$ 1.6 31.5 $\pm$ 3.6 40.2 $\pm$ 5.1 14.6 $\pm$ 2.7 16.0 $\pm$ 3.7 25.9 $\pm$ 2.2 8.9 $\pm$ 1.2 8.4 $\pm$ 2.2 8.9 $\pm$ 2.7
29.4
SPLADE++ 66.8 $\pm$ 1.3 8.2 $\pm$ 1.2 34.0 $\pm$ 3.4 47.9 $\pm$ 5.0 9.5 $\pm$ 2.2 8.9 $\pm$ 2.9 27.5 $\pm$ 2.2 9.5 $\pm$ 1.2 11.4 $\pm$ 2.2 13.9 $\pm$ 3.3
29.8
SPLADE-v2 67.9 $\pm$ 1.4 10.6 $\pm$ 2.1 37.8 $\pm$ 3.7 40.7 $\pm$ 5.0 16.2 $\pm$ 3.0 20.2 $\pm$ 4.1 22.3 $\pm$ 2.3 7.2 $\pm$ 1.1 10.7 $\pm$ 2.1 11.0 $\pm$ 3.0
31.0
SPECTER-v2 71.4 $\pm$ 1.2 13.5 $\pm$ 2.3 37.7 $\pm$ 3.3 49.1 $\pm$ 5.1 9.8 $\pm$ 2.0 13.0 $\pm$ 3.4 30.4 $\pm$ 2.1 11.2 $\pm$ 1.2 7.7 $\pm$ 1.6 12.0 $\pm$ 3.1
31.4
E5-L-v2 72.0 $\pm$ 1.2 14.6 $\pm$ 1.9 43.5 $\pm$ 3.2 62.0 $\pm$ 4.8 36.6 $\pm$ 4.0 39.9 $\pm$ 4.8 29.4 $\pm$ 2.7 10.5 $\pm$ 1.7 11.1 $\pm$ 2.6 14.9 $\pm$ 3.4
38.5
SIMCSE 70.8 $\pm$ 1.3 14.4 $\pm$ 1.8 39.8 $\pm$ 3.5 51.1 $\pm$ 5.1 31.7 $\pm$ 4.0 37.0 $\pm$ 5.0 27.7 $\pm$ 2.7 10.5 $\pm$ 1.6 12.9 $\pm$ 2.5 15.6 $\pm$ 3.7
36.6
Promptagator 76.0 $\pm$ 1.4 15.6 $\pm$ 1.8 63.0 $\pm$ 3.3 77.8 $\pm$ 4.2 59.1 $\pm$ 4.2 62.9 $\pm$ 4.7 NA NA NA NA
NA
Encoder-Decoder Models
TART 64.9 $\pm$ 1.5 10.2 $\pm$ 1.7 47.6 $\pm$ 3.1 63.7 $\pm$ 4.8 42.7 $\pm$ 3.7 46.1 $\pm$ 4.9 32.8 $\pm$ 2.5 10.8 $\pm$ 1.3 9.4 $\pm$ 1.9 7.9 $\pm$ 2.7
39.5
TART+O 49.6 $\pm$ 1.8 4.9 $\pm$ 1.0 23.1 $\pm$ 2.8 28.0 $\pm$ 4.4 17.9 $\pm$ 2.4 20.7 $\pm$ 4.1 23.6 $\pm$ 2.2 7.7 $\pm$ 1.0 14.3 $\pm$ 2.3 18.3 $\pm$ 3.8
25.7
MonoT5 66.9 $\pm$ 1.5 12.4 $\pm$ 1.8 45.9 $\pm$ 3.2 52.0 $\pm$ 4.9 50.8 $\pm$ 4.2 59.0 $\pm$ 4.9 33.2 $\pm$ 2.5 14.2 $\pm$ 2.3 13.9 $\pm$ 2.7 11.9 $\pm$ 3.2
42.1
Decoder-Only Model
RankLLaMA 75.1 $\pm$ 1.3 18.4 $\pm$ 2.0 53.2 $\pm$ 2.9 71.9 $\pm$ 4.4 63.7 $\pm$ 3.9 69.8 $\pm$ 4.6 29.7 $\pm$ 2.2 9.8 $\pm$ 1.1 15.4 $\pm$ 2.4 19.0 $\pm$ 4.0
47.4
@@
RankStripedHyena 71.7 $\pm$ 1.3 14.6 $\pm$ 1.9 40.7 $\pm$ 3.1 59.8 $\pm$ 5.0 30.4 $\pm$ 3.5 37.9 $\pm$ 4.8 30.5 $\pm$ 2.6 10.6 $\pm$ 1.7 10.6 $\pm$ 2.4 14.9 $\pm$ 3.4
36.8
Rank+OStripedHyena 70.6 $\pm$ 1.3 15.5 $\pm$ 1.9 41.4 $\pm$ 3.4 48.9 $\pm$ 5.2 34.2 $\pm$ 3.9 40.9 $\pm$ 4.9 28.7 $\pm$ 2.4 10.1 $\pm$ 1.2 11.3 $\pm$ 2.5 14.9 $\pm$ 3.4
37.2
RankLlama2-7b 71.0 $\pm$ 1.3 13.2 $\pm$ 1.6 41.4 $\pm$ 2.8 54.8 $\pm$ 5.1 29.5 $\pm$ 3.3 37.8 $\pm$ 4.8 29.3 $\pm$ 2.6 37.8 $\pm$ 1.8 10.5 $\pm$ 2.3 11.8 $\pm$ 3.1
36.3
Rank+OLlama2-7b 70.8 $\pm$ 1.3 14.1 $\pm$ 1.7 41.1 $\pm$ 3.0 54.6 $\pm$ 4.9 26.7 $\pm$ 3.0 34.8 $\pm$ 4.6 28.8 $\pm$ 2.4 9.7 $\pm$ 1.3 10.6 $\pm$ 2.2 10.8 $\pm$ 3.0
35.6
RankLlama2-13b 71.2 $\pm$ 1.4 13.0 $\pm$ 1.6 40.1 $\pm$ 2.9 53.9 $\pm$ 5.1 26.5 $\pm$ 3.1 36.8 $\pm$ 4.8 29.6 $\pm$ 2.7 10.6 $\pm$ 1.5 10.7 $\pm$ 2.3 14.9 $\pm$ 3.5
35.6
Rank+OLlama2-13b 70.5 $\pm$ 1.3 14.2 $\pm$ 1.7 39.1 $\pm$ 2.9 51.7 $\pm$ 5.2 28.5 $\pm$ 3.3 35.8 $\pm$ 4.7 28.4 $\pm$ 2.7 10.0 $\pm$ 1.6 8.9 $\pm$ 2.0 12.8 $\pm$ 3.2
35.0
RankLlama2-70b 70.7 $\pm$ 1.3 15.1 $\pm$ 1.9 39.0 $\pm$ 2.8 49.8 $\pm$ 5.0 27.6 $\pm$ 3.2 36.8 $\pm$ 4.7 30.2 $\pm$ 2.7 10.3 $\pm$ 1.4 10.5 $\pm$ 2.3 13.8 $\pm$ 3.4
35.6
Rank+OLlama2-70b 70.4 $\pm$ 1.3 14.2$\pm$ 1.7 41.6$\pm$ 3.0 51.7 $\pm$ 5.0 27.6 $\pm$ 3.1 36.8$\pm$ 4.7 28.8 $\pm$ 2.4 9.7 $\pm$ 1.3 10.5 $\pm$ 2.4 13.9$\pm$ 3.3
35.8
RankGPT4 76.2 $\pm$ 1.3 17.5 $\pm$ 2.3 26.7 $\pm$ 2.2 17.0 $\pm$ 3.8 79.5 $\pm$ 3.5 85.9 $\pm$ 3.4 39.9 $\pm$ 2.6 14.8 $\pm$ 1.4 40.1 $\pm$ 4.0 48.0 $\pm$ 5.1
52.5
Rank+OGPT4 77.4 $\pm$ 1.3 17.6 $\pm$ 2.1 54.4 $\pm$ 3.2 70.9 $\pm$ 4.5 82.1 $\pm$ 3.3 88.0 $\pm$ 3.2 38.6 $\pm$ 2.8 14.6 $\pm$ 1.5 62.3 $\pm$ 3.9 70.1 $\pm$ 4.5
63.0

### v2

@@ 3 A Benchmark for LLM-based IR Systems
Num. of Triplets Num. of Queries Num. of Docs Relevant Doc. per Query Query Length Doc Length Lexical Overlap
@@
DORIS-MAE AI/ML Retrieve relevant abstracts for the research question. Multi-level 3k 60 110.6 18.2 166.7 199.5 0.147
ArguAna Debate Retrieve the most effective argument that counters the query. Binary 19k 100 50.0 1 189.1 185.5 0.139
WhatsThatBook Literature Find the book from user’s ambiguous descriptions. Binary 3k 100 50.4 1 152.2 189.3 0.049
Clinical-Trial BioMed Match patient records with suitable clinical trials. 3-level 1k 50 68.4 19.7 78.5 175.6 0.050
RELIC Literature Recover the masked quotation in a literary analysis. Binary 68k 100 50.6 1 169.1 85.4 0.027
BIRCO Avg. - - - 20k 51.3 66.0 8.4 168.3 185.1 0.083
BEIR Avg. - - - - 3710.9 2417k 43.5 9.8 124.9 0.384
@@
Max possible 100.0 100.0 Max possible 100.0 94.3 Max possible 100.0 100.0
ANCEFirstP 33.0 20.0 ANCEFirstP 81.9 49.6 E5-L-v2 50.0 38.5
SimLM 41.1 18.1 SimLM 85.2 44.6 RankLLaMA 56.6 47.4
SPLADE-v2 36.8 17.7 BM25 59.1 33.5 TART 44.8 39.5
@@ 3.3.2 Reducing contamination
MS MARCO 76% 18% 6%
TREC-COVID 94% 4% 2%
NQ 89% 7% 4%
HotpotQA 60% 20% 20%
FiQA 73% 19% 8%
@@
DBPedia 78% 12% 10%
FEVER 99% 0% 1%
Climate-FEVER 81% 18% 1%
Touché-2020 98% 2% 0%
SciFact 71% 25% 4%
Table 3: Data contamination in BEIR. GPT4 answers the query without access to any candidate passages. C0,1,2 corresponds to Category 0,1,2. C0 means contamination. Numbers are rounded to the nearest percentile.
@@ 3.8 Candidate Pool Construction
We construct candidate pools for each query using the lexicon-matching algorithm BM25 Robertson and Walker (1994); Trotman et al. (2014) and state-of-the-art text embedding model ada-002 Greene et al. (2022). Each query has a candidate pool of 50 passages. (The ground-truth passage is inserted when necessary.) As shown in Appendix C and Table 9, the difficulty is comparable to the original datasets for four of five tasks, and remains challenging for the fifth.
@@ 5 LLM Benchmark Results
nDCG@10 Recall@5 nDCG@10 Recall@5 nDCG@10 Recall@5 nDCG@10 Recall@5 nDCG@10 Recall@5 Avg. nDCG@10
Maximum Possible 100.0 45.6 100.0 100.0 100.0 100.0 100.0 43.0 100.0 100.0 100.0
Random 49.9 4.5 9.1 9.9 8.9 9.8 22.4 8.4 8.9 10.0 19.8
@@
TART 64.9 10.2 49.3 66.7 42.7 46.1 32.9 11.1 9.1 7.0 39.8
TART+O 49.6 4.9 23.6 29.9 17.9 20.7 24.1 7.5 13.2 17.1 25.7
MonoT5 66.9 12.4 46.8 52.7 50.8 59.0 34.3 15.6 11.9 10.0 42.1
@@
RankStripedHyena 71.7 14.6 39.8 58.0 33.6 39.8 29.5 10.3 10.9 13.0 37.1
Rank+OStripedHyena 70.6 15.5 38.9 54.8 32.8 39.9 29.8 10.8 9.2 11.9 36.3
RankLlama2-7b 71.0 13.2 40.5 58.9 25.7 35.9 31.2 11.5 11.1 11.9 35.9
Rank+OLlama2-7b 70.8 14.1 39.4 50.9 26.5 37.8 28.5 9.6 10.8 14.0 35.2
RankLlama2-13b 71.2 13.0 41.1 57.0 27.1 37.8 32.9 13.8 9.9 14.9 36.4
Rank+OLlama2-13b 70.5 14.2 37.8 47.8 28.0 38.9 30.0 10.9 9.2 12.9 35.1
RankLlama2-70b 70.7 15.1 41.3 55.8 30.2 37.8 29.0 9.8 10.6 12.9 36.4
Rank+OLlama2-70b 70.4 14.2 40.2 55.9 25.6 38.9 30.0 11.4 9.1 10.9 35.1
RankGPT4 76.2 17.5 27.3 17.0 80.1 86.9 40.3 17.0 40.6 49.3 52.9
Rank+OGPT4 77.4 17.7 55.6 70.0 82.1 88.0 38.6 17.0 63.4 71.1 63.4
@@
Score+OStripedHyena 67.5 9.5 21.6 23.1 43.0 50.8 31.8 10.8 21.1 24.8 37.0
Reason+OStripedHyena 65.9 10.9 27.7 32.0 23.1 23.1 29.2 13.8 12.7 11.9 31.7
Subtask+OStripedHyena 74.8 15.5 16.3 20.0 43.5 53.0 32.3 10.9 24.7 29.1 38.3
Score+OLlama2-7b 71.6 14.1 29.3 40.1 30.5 34.7 31.1 11.2 10.9 14.0 34.7
Reason+OLlama2-7b 72.1 13.8 21.9 27.0 21.4 28.1 29.9 9.7 10.1 10.0 31.1
Subtask+OLlama2-7b 69.2 11.8 22.3 27.9 29.2 31.1 28.7 10.0 9.4 11.9 31.8
Score+OLlama2-13b 77.3 14.5 15.0 24.1 29.6 34.0 30.6 10.5 16.2 18.0 33.7
Reason+OLlama2-13b 60.1 7.5 32.6 45.1 19.4 28.0 25.4 8.7 9.1 8.9 29.3
Subtask+OLlama2-13b 71.3 14.1 18.1 20.8 52.8 61.9 27.7 10.7 19.7 20.0 37.9
Score+OLlama2-70b 75.1 15.1 41.6 53.0 60.8 70.0 36.3 12.6 11.7 14.9 45.1
Reason+OLlama2-70b 70.1 14.5 45.0 60.1 43.4 48.0 32.6 11.6 12.1 15.9 40.6
Subtask+OLlama2-70b 76.7 16.3 21.0 25.0 42.6 50.2 33.3 11.8 17.3 21.0 38.2
Score+OGPT4 79.9 19.3 53.2 70.8 83.3 90.9 43.7 19.6 56.9 73.2 63.4
Reason+OGPT4 74.9 18.1 61.5 76.1 75.2 82.8 44.6 16.4 41.7 52.8 59.6
Subtask+OGPT4 78.5 18.5 70.8 83.9 77.9 85.9 44.3 19.5 53.3 66.0 65.0
@@ Appendix A Experiment Details
nDCG@10 Recall@5 nDCG@10 Recall@5 nDCG@10 Recall@5 nDCG@10 Recall@5 nDCG@10 Recall@5 Avg. nDCG@10
Maximum Possible 100.0 45.6 100.0 100.0 100.0 100.0 100.0 43.0 100.0 100.0 100.0
Random 49.9 4.5 9.1 9.9 8.9 9.8 22.4 8.4 8.9 10.0 19.8
@@
RoBERTa-L 66.8 $\pm$ 1.3 12.0 $\pm$ 1.6 31.8 $\pm$ 3.5 41.2 $\pm$ 4.8 14.6 $\pm$ 2.7 16.0 $\pm$ 3.7 26.6 $\pm$ 2.3 9.4 $\pm$ 1.4 7.0 $\pm$ 1.8 7.8 $\pm$ 2.6 29.4
SPLADE++ 66.8 $\pm$ 1.3 8.2 $\pm$ 1.2 34.2 $\pm$ 3.3 47.6 $\pm$ 4.9 9.5 $\pm$ 2.2 8.9 $\pm$ 2.9 28.0 $\pm$ 2.2 10.8 $\pm$ 1.5 10.4 $\pm$ 2.1 12.9 $\pm$ 3.4 29.8
SPLADE-v2 67.9 $\pm$ 1.4 10.6 $\pm$ 2.1 38.5 $\pm$ 3.6 41.9 $\pm$ 4.8 16.2 $\pm$ 3.0 20.2 $\pm$ 4.1 23.0 $\pm$ 2.3 7.5 $\pm$ 1.2 10.4 $\pm$ 2.2 11.0 $\pm$ 3.1 31.2
E5-L-v2 72.0 $\pm$ 1.2 14.6 $\pm$ 1.9 44.9 $\pm$ 3.2 62.0 $\pm$ 4.8 36.6 $\pm$ 4.0 39.9 $\pm$ 4.8 30.4 $\pm$ 2.6 11.3 $\pm$ 1.8 10.9 $\pm$ 2.5 14.9 $\pm$ 3.6 39.0
SIMCSE 70.8 $\pm$ 1.3 14.4 $\pm$ 1.8 40.3 $\pm$ 3.5 51.9 $\pm$ 5.1 31.7 $\pm$ 4.0 37.0 $\pm$ 5.0 27.1 $\pm$ 2.8 10.7 $\pm$ 1.6 12.7 $\pm$ 2.4 15.9 $\pm$ 3.6 36.5
Promptagator 76.0 $\pm$ 1.4 15.6 $\pm$ 1.8 60.9 $\pm$3.2 73.7 $\pm$ 4.4 58.4 $\pm$ 4.2 62.9 $\pm$ 4.7 NA NA NA NA NA
Finetuned Encoder-Decoder Models
TART 64.9 $\pm$ 1.5 10.2 $\pm$ 1.7 49.3 $\pm$ 3.4 66.7 $\pm$ 4.8 42.7 $\pm$ 3.7 46.1 $\pm$ 4.9 32.9 $\pm$ 2.7 11.1 $\pm$ 1.4 9.1 $\pm$ 1.7 7.0 $\pm$ 2.5 39.8
TART+O 49.6 $\pm$ 1.8 4.9 $\pm$ 1.0 23.6 $\pm$ 2.8 29.9 $\pm$ 4.5 17.9 $\pm$ 2.4 20.7 $\pm$ 4.1 24.1 $\pm$ 2.3 7.5 $\pm$ 1.1 13.2 $\pm$ 2.2 17.1 $\pm$ 3.7 25.7
MonoT5 66.9 $\pm$ 1.5 12.4 $\pm$ 1.8 46.8 $\pm$ 3.4 52.7 $\pm$ 5.1 50.8 $\pm$ 4.2 59.0 $\pm$ 4.9 34.3 $\pm$ 2.5 15.6 $\pm$ 2.5 11.9 $\pm$ 2.5 10.0 $\pm$ 3.0 42.1
Finetuned Decoder-Only Model
RankLLaMA 75.1 $\pm$ 1.3 18.4 $\pm$ 2.0 54.1 $\pm$ 3.0 72.6 $\pm$ 4.5 63.7 $\pm$ 3.9 69.8 $\pm$ 4.6 30.7 $\pm$ 2.2 10.9 $\pm$ 1.4 14.9 $\pm$ 2.4 18.0 $\pm$ 3.7 47.7
@@
RankSH 71.7 $\pm$ 1.3 14.6 $\pm$ 1.8 39.8 $\pm$ 3.0 58.0 $\pm$ 4.9 33.6 $\pm$ 3.8 39.8 $\pm$ 4.8 29.5 $\pm$ 2.2 10.3 $\pm$ 1.5 10.9 $\pm$ 2.5 13.0 $\pm$ 3.3 37.1
Rank+OSH 70.6 $\pm$ 1.3 15.5 $\pm$ 1.9 38.9 $\pm$ 3.0 54.8 $\pm$ 5.0 32.8 $\pm$ 3.8 39.9 $\pm$ 4.8 29.8 $\pm$ 2.4 10.8 $\pm$ 1.5 9.2 $\pm$ 2.0 11.9 $\pm$ 3.2 36.3
RankLM2-7b 71.0 $\pm$ 1.3 13.2 $\pm$ 1.6 40.5 $\pm$ 3.1 58.9 $\pm$ 4.8 25.7 $\pm$ 2.8 35.9 $\pm$ 4.7 31.2 $\pm$ 2.6 11.5 $\pm$ 1.6 11.1 $\pm$ 2.5 11.9 $\pm$ 3.2 35.9
Rank+OLM2-7b 70.8 $\pm$ 1.3 14.1 $\pm$ 1.7 39.4 $\pm$ 2.9 50.9 $\pm$ 4.9 26.5 $\pm$ 2.9 37.8 $\pm$ 4.8 28.5 $\pm$ 2.5 9.6 $\pm$ 1.5 10.8 $\pm$ 2.4 14.0 $\pm$ 3.5 35.2
RankLM2-13b 71.2 $\pm$ 1.4 13.0 $\pm$ 1.6 41.1 $\pm$ 2.9 57.0 $\pm$ 4.8 27.1 $\pm$ 3.0 37.8 $\pm$ 4.9 32.9 $\pm$ 3.0 13.8 $\pm$ 2.3 9.9 $\pm$ 2.0 14.9 $\pm$ 3.6 36.4
Rank+OLM2-13b 70.5 $\pm$ 1.3 14.2 $\pm$ 1.7 37.8 $\pm$ 2.9 47.8 $\pm$ 5.0 28.0 $\pm$ 3.1 38.9 $\pm$ 4.9 30.0 $\pm$ 2.5 10.9 $\pm$ 1.5 9.2 $\pm$ 2.1 12.9 $\pm$ 3.3 35.1
RankLM2-70b 70.7 $\pm$ 1.3 15.1 $\pm$ 1.9 41.3 $\pm$ 3.1 55.8 $\pm$ 4.9 30.2 $\pm$ 3.3 37.8 $\pm$ 4.8 29.0 $\pm$ 2.5 9.8 $\pm$ 1.5 10.6 $\pm$ 2.4 12.9 $\pm$ 3.4 36.4
Rank+OLM2-70b 70.4 $\pm$ 1.3 14.2 $\pm$ 1.7 40.2 $\pm$ 3.1 55.9 $\pm$ 5.0 25.6 $\pm$ 2.9 38.9 $\pm$ 4.8 30.0 $\pm$ 2.5 11.4 $\pm$ 1.8 9.1 $\pm$ 2.0 10.9 $\pm$ 3.1 35.1
RankGPT4 76.2 $\pm$ 1.3 17.5 $\pm$ 2.3 27.3 $\pm$ 2.2 17.0 $\pm$ 3.7 80.1 $\pm$ 3.4 86.9 $\pm$ 3.3 40.3 $\pm$ 2.7 17.0 $\pm$ 2.4 40.6 $\pm$ 3.9 49.3 $\pm$ 4.8 52.9
Rank+OGPT4 77.4 $\pm$ 1.3 17.7 $\pm$ 2.1 55.6 $\pm$ 3.4 70.0 $\pm$ 4.7 82.1 $\pm$ 3.3 88.0 $\pm$ 3.2 38.6 $\pm$ 2.9 17.0 $\pm$ 2.5 63.4 $\pm$ 4.0 71.1 $\pm$ 4.6 63.4